If one side of a square is increased by 8 cm and an adjacent side decreased by 2 cm a rectangle is formed whose permeter is 40 cm. Find the length of a side of the square.
Answer by mananth(16946)

Let the length of the side of square be x
x+8, x-2 are the sides of the rectangle.
Perimeter = 2L+2W
2(x-2)+2(x+8)=40
2x-4+2x+16=40
4x+12=404x=28
x=7 cm the length of side of square